# RiteWeight: Randomized Iterative Trajectory Reweighting for Steady-State Distributions Without Discretization Error

### v3

@@ Abstract
A significant challenge in molecular dynamics (MD) simulations is ensuring that sampled configurations converge to the equilibrium or nonequilibrium stationary distribution of interest. Lack of convergence constrains the estimation of free energies as well as rates and mechanisms of molecular transitions. Here, we introduce the “Randomized ITErative trajectory reWeighting” (RiteWeight) algorithm to estimate a stationary distribution from unconverged simulation data. This method iteratively reweights trajectory segments in a self-consistent way by solving for the stationary distribution of a Markov state model (MSM), updating segment weights, and employing a new random clustering in each iteration. The repeated clustering mitigates the phase-space discretization error inherent in existing trajectory reweighting techniques and yields quasi-continuous configuration-space distributions. RiteWeight accurately recovers the stationary distribution, even without requiring the Markov property at the cluster level. We present mathematical analysis of the RiteWeight fixed point. We empirically validate the method using both synthetic MD Trp-cage trajectories, for which the stationary solution is exactly calculable, and standard atomistic MD Trp-cage trajectories extracted from a long reference simulation. In both test systems, RiteWeight corrects flawed distributions and generates accurate observables for equilibrium and nonequilibrium steady states. The results highlight the value of correcting the underlying trajectory distribution rather than using a standard MSM.
@@
Molecular dynamics (MD) simulation is a key tool for studying the behavior of proteins and other biomolecules, but despite four decades of hardware and algorithm advances, MD cannot completely characterize the biomolecular behavior of most systems of interest. Typically, the molecular configurations generated by MD and related methods fail to conform to the equilibrium or nonequilibrium steady-state distribution of interest, thereby limiting the accuracy of computed observables such as rate constants. The present report introduces a novel approach for correcting mis-distributed configurations using a Randomized ITErative reWeighting (RiteWeight) strategy. The approach is validated for protein folding systems under both equlibrium and nonequilibrium conditions.
@@ 1 Introduction
Despite advances in enhanced equilibrium and path sampling methods [1, 2, 3, 4, 5, 6, 7], the study of complex biomolecular systems remains resource-intensive, often surpassing researchers’ computational capabilities. Conventional molecular dynamics (MD) cannot produce well-sampled conformational distributions except in special cases [8]. A major challenge is estimating the “steady state” or stationary distribution of the sampled states [9, 10, 11]. The accuracy of the stationary distribution is crucial because it reveals thermodynamic and kinetic properties of the system. The stationary distribution is needed to calculate free energies as well as identifying mechanistic pathways, rate constants, and committor reaction coordinates [12, 13, 14].
@@
Recently, methods based on AlphaFold have generated protein structural ensembles [15, 16, 17, 18]. However, these methods are heuristic in the sense that they are not designed to produce equilibrium ensembles conforming to the Boltzmann factor. The algorithm described here, however, can leverage heuristic ensembles as the starting point for conventional MD and subsequent Boltzmann-factor reweighting. While heuristic starting configurations can also be updated with enhanced sampling [19], all such approaches can be improved by fixing the sample weights.
@@
One principled way to estimate the steady-state distribution from sampled trajectory data is to build a “Markov state model” (MSM), i.e., a discrete-state transition matrix that leads to an approximation of the stationary distribution [20, 21]. However, previous work has noted that MSM estimates of stationarity are biased by the trajectory training data [22]. To ameliorate this issue, a further step can be taken in a MSM pipeline, namely, the trajectories can be reweighted based on the matrix’s stationary solution [13]. This procedure — referred to here as “single-shot reweighting” — can be helpful. However, it cannot correct the weights of trajectories within the chosen discrete states, nor can it provide a set of weights consistent with a subsequently computed transition matrix [12, 23], potentially skewing the estimation of observables.
@@
Unbiased estimation only occurs in the traditional MSM framework when the trajectories in each discrete state are locally consistent with the stationary distribution for the chosen boundary conditions. For example, source-sink boundary conditions lead to the challenging requirement to sample from the nonequilibrium steady state [12, 13, 24]. The need for unbiased equilibrium or nonequilibrium samples creates a chicken and egg problem which can be addressed by an iterative solution [12, 23], at the cost of generating long trajectories. We note that non-traditional MSMs can be used to obtain unbiased estimates of observables [14], but traditional MSMs are biased for both equilibrium and nonequilibrium observables even when significant trajectory data is available [22, 25].
@@
As an approach for correcting standard MSMs, we here introduce the “Randomized ITErative trajectory reWeighting” (RiteWeight) algorithm. RiteWeight reweights trajectories generated without biasing forces into their correct stationary distribution when sufficient data is available. The method applies to both equilibrium and nonequilibrium steady states. It naturally uses trajectories of any length, as short as just a single time step, as it does not rely on dynamical relaxation. Thus, RiteWeight can employ data generated from standard molecular dynamics or path sampling approaches, so long as no biasing forces have been used. We emphasize that RiteWeight is distinct from conventional importance sampling, which requires a known and well-sampled initial distribution [26, 27].
@@
Figure 1 shows how the RiteWeight algorithm iteratively reweights trajectories using the stationary measure for a discrete-state transition matrix , which is determined in each iteration by a new random clustering. The relative weights of the trajectory segments in each cluster are fixed during a single iteration. In the equilibrium case, the transition matrix used is akin to a MSM based on weighted trajectories. In the nonequilibrium case, the matrix is computed from weighted trajectories conforming to source-sink boundary conditions, akin to a “history augmented” MSM [28, 29, 25, 24]. However, RiteWeight uses no history information and thus eliminates the requirement that all segments to be traced back to the source state. RiteWeight also does not require the Markov property to hold at the cluster level, because discrete states are used only to estimate stationarity and not to propagate dynamics over time.
@@
RiteWeight differs from prior strategies for reweighting sampled trajectory data. For instance, one approach explicitly approximates the configuration-space density function in order to reweight configurations [30]. Another method updates the weights using a variational principle for the equilibrium distribution [31]. RiteWeight, however, does not require density estimation and only uses standard Markov models without additional fitting parameters or assumed functional forms. In prior work, our research group has applied iterative reweighting without changing cluster definitions [23], but this approach does not yield the quasi-continuous distribution which is achieved using RiteWeight.
@@
The present report evaluates the RiteWeight algorithm through mathematical analysis and empirical testing. First, we present mathematical analysis identifying the fixed point of the RiteWeight algorithm. Then, we present numerical tests based on both synthetic molecular dynamics (SynMD) and true MD trajectories of the Trp-cage miniprotein. SynMD consists of trajectories generated from a fine-grained MSM with each state mapped to an atomistic conformation, and it enables comparison to an exactly calculable reference distribution [32]. The true MD data is a single 208 s trajectory generated by the Shaw group [8]. We compare the performance of RiteWeight, single-shot reweighting, and traditional MSMs for estimating the stationary distribution from mis-distributed data sets. Overall, RiteWeight yields better agreement with reference values for all observables considered in both equilibrium and nonequilibrium scenarios. RiteWeight achieves good performance using extremely short trajectory segments, enabling computation of mechanistic, path-based observables. Last, RiteWeight results are independent of the number of clusters used to discretize the phase space.
@@ 2 RiteWeight Algorithm
The RiteWeight algorithm estimates steady-state probability distributions using kinetic information extracted from numerous short MD trajectories or one long trajectory. The algorithm processes sequential pairs of phase-space points generated by unbiased dynamics and separated by a fixed lag time: . We refer to these pairs as “trajectory segments”. Unlike MSMs, the algorithm’s ability to estimate stationary probabilities is not constrained by the choice of lag time [21, 20]. Unlike single-shot reweighting, which requires within-cluster stationarity for unbiased results [12, 13, 24], RiteWeight imposes no such constraints. Therefore, transition pair data can be collected right from the start of the trajectories, ensuring that no data is discarded and all available information is leveraged in the analysis.
@@ 2.1 The algorithm
Assign an initial weight to each trajectory segment connecting two consecutive configurations, using the normalization . A possible choice is uniform weights , where is the number of segments.
@@
Calculate the stationary probabilities for the discrete states using the left leading eigenvector of the matrix , i.e., .
@@
(2) where is the previous weight of the trajectory segment, is the new weight, is the current iteration’s estimate of the stationary probability for the cluster , and .
@@
Several points are noteworthy. In step 2, there are two approaches for assigning the initial weights to trajectory segments. One option is to assign uniform weights to all segments, which serves as an uninformative prior. Alternatively, if there is prior knowledge regarding the stationary distribution of the system, this information can be used to assign the initial weights in a more informed manner. Random clusters are defined in step 3, and a new clustering is introduced at every iteration. In step 6, the weight of each trajectory in cluster is updated to be the currently estimated stationary probability multiplied by a fraction , where is the previous estimate for the trajectory segment and is the previous estimate for the cluster.
Last, we define convergence in the numerical experiments by running RiteWeight for – iterations, until the weights or mean first passage time estimate exhibits small, steady fluctuations. Then, we average the weights over the final 1,000 or 10,000 iterations to further reduce noise. See supplementary Figs. 7, 8, and 10 for examples of RiteWeight convergence analysis.
@@ 2.2 Analysis of RiteWeight fixed point
For simplicity and clarity, our analysis assumes that the configurations all belong to a discrete state space of “microstates” indexed by or . We expect that RiteWeight in continuous space behaves similarly to RiteWeight in discrete space at a level of resolution automatically determined by the algorithm and its stopping criterion.
@@
Here, is the initial weight assigned to segment . The Appendix proves the stationary vector of defines the fixed point of RiteWeight assuming this stationary vector is unique and the clusters are chosen according to a random hyperplane tesselation.
The fixed point analysis suggests that RiteWeight will converge to the true stationary distribution given sufficiently dense and unbiased local sampling. In the limit of high transition counts and small microstates, the matrix approaches the true transition operator. Thus, RiteWeight will find the correct distribution regardless of the number of clusters or the lag time. However, in more realistic settings, the accuracy of RiteWeight is constrained by the data analyzed, pointing to a valuable role for the adaptive sampling methods that will be mentioned in the Discussion section.
Given the description of the RiteWeight fixed point, it is natural to ask whether the RiteWeight stationary distribution can be derived directly from a fine-grained MSM transition matrix. In practice, this is not possible because the resolution of the microstates corresponding to RiteWeight cannot be easily determined for an arbitrary data set. The Results section will show the limitations of simply using a fine-grained Markov state model without RiteWeight.
@@ 3 Test systems
Here we introduce the synthetic MD and atomistic MD systems that we used in our empirical RiteWeight experiments.
@@ 3.1 Synthetic MD for Trp-cage
Model: The SynMD model for the Trp-cage miniprotein [32] is derived from an MSM trained on a 208 s atomistic MD trajectory [8]. The MSM uses 10,500 states, which is a finer discretization than a typical MSM, and each state is mapped to a specific atomistic configuration. Kinetic Monte Carlo then yields a “synthetic” MD trajectory that closely matches the statistical behavior of the original atomistic MD [32]. The MSM lag time and hence the interval between synMD configurations is 1 ns, and the mapped configurations can be processed using standard MSMs or RiteWeight. Because the synMD model is governed by an explicitly known MSM transition matrix, we can calculate the equilibrium distribution explicitly using the leading left eigenvector of the matrix.
Trajectory preparation: We prepared a data set of 5 ns SynMD trajectories using a sampling measure far from equilibrium. First we applied tICA with a 5 ns lag time [25] to the set of minimal residue-residue distances, which are defined by the closest distances between the heavy atoms of two residues separated in sequence by at least two neighboring residues. Then we assigned indices to the 10,500 MSM states based on the slowest time-lagged independent component (TIC1) [32]. We partitioned the SynMD indices into 500 bins and selected 20 initial states randomly with replacement from each bin. For each selected state, we initiated a short trajectory of 5 ns, i.e., 5 steps. Hence the unprocessed trajectory data consists of 10,000 trajectory segments, each 5 steps long, distributed roughly evenly along the TIC1 coordinate.
RiteWeight Analysis: We applied RiteWeight in a manner blind to the discrete nature of synMD, using only the mapped atomistic configurations. We defined features using the slowest 10 tICs, and we rescaled the tICs based on commute distances. Then we performed RiteWeight using either or clusters based on the rescaled tICs.
@@ 3.2.1 Equilibrium distribution
Trajectory preparation: Again we prepared a data set of short trajectories using a sampling measure far from equilibrium. First we applied tICA with a 10 ns lag time [25] to the set of minimal residue-residue distances. Then we projected all conformations from the 208 s trajectory onto TIC1 and grouped them into 100 uniformly spaced bins. We subsampled 10,000 conformations within each bin; however, if fewer than 10,000 conformations were available in a bin, we selected all the conformations. We used these conformations as the starting points for segments consisting of two snapshots from the long trajectory separated by a 10 ns lag time.
RiteWeight and MSM Analysis: We retained a sufficient number of tICs to explain 95% of the total variance, and we rescaled the tICs based on commute distances. For RiteWeight, we performed random clustering () using the rescaled tICs as input features. For the MSM analysis, we performed -means clustering based on – clusters. For all downstream MSM analysis following clustering, we used the PyEMMA software package [34].

### v4

@@ Abstract
A significant challenge in molecular dynamics (MD) simulations is ensuring that sampled configurations converge to the equilibrium or nonequilibrium stationary distribution of interest. Lack of convergence constrains the estimation of free energies and of rates and mechanisms for molecular transitions. Here, we introduce the “Randomized ITErative trajectory reWeighting” (RiteWeight) algorithm to estimate a stationary distribution from unconverged simulation data. This method iteratively reweights trajectory segments in a self-consistent way by solving for the stationary distribution of a Markov state model (MSM), updating segment weights, and employing a new random clustering in each iteration. The repeated clustering mitigates the configuration-space discretization error inherent in existing trajectory reweighting techniques and yields quasi-continuous configuration-space distributions. RiteWeight accurately recovers the stationary distribution even without requiring the Markov property at the cluster level. We present mathematical analysis of the RiteWeight fixed point. We empirically validate the method using both synthetic MD Trp-cage trajectories, for which the stationary solution is exactly calculable, and standard atomistic MD Trp-cage trajectories, which are extracted from a long reference simulation. In both test systems, RiteWeight corrects flawed distributions and generates accurate observables for equilibrium and nonequilibrium steady states. The results highlight the value of correcting the underlying trajectory distribution rather than using a standard MSM.
@@
Molecular dynamics (MD) simulation is a key tool for studying the behavior of proteins and other biomolecules, but despite four decades of hardware and algorithm advances, MD cannot characterize the biomolecular behavior of most systems of interest. Typically, the molecular configurations generated by MD and related methods fail to conform to the equilibrium or nonequilibrium steady state distribution of interest, thereby limiting the accuracy of computed observables such as rate constants. The present report introduces a novel approach for correcting mis-distributed configurations using a Randomized ITErative reWeighting (RiteWeight) strategy. The approach is validated for protein folding systems under both equilibrium and nonequilibrium conditions.
@@ 1 Introduction
Despite advances in enhanced equilibrium and path sampling methods [6, 14, 27, 15, 24, 48, 17], the study of complex biomolecular systems remains resource-intensive, often surpassing researchers’ computational capabilities. Conventional molecular dynamics (MD) cannot produce well-sampled configurational distributions except in special cases [26]. A major challenge is estimating the “steady state” or stationary distribution of the sampled configurations [49, 43, 7]. The accuracy of the stationary distribution is crucial because it reveals thermodynamic and kinetic properties of the system. The stationary distribution is needed to calculate free energies as well as identify mechanistic pathways, rate constants, and committor reaction coordinates [33, 46, 28].
@@
Recently, methods based on AlphaFold have generated protein structural ensembles [21, 31, 22, 37]. However, these methods are heuristic in the sense that they are not designed to produce equilibrium ensembles conforming to the Boltzmann factor. The algorithm described here, however, can leverage heuristic ensembles as the starting point for conventional MD and subsequent Boltzmann-factor reweighting. While heuristic starting configurations can be updated with enhanced sampling [45], all such approaches can be improved by correcting the sample weights.
@@
One principled way to estimate the steady state distribution from sampled trajectory data is to build a “Markov state model” (MSM), i.e., a discrete-state transition matrix that leads to an approximation of the stationary distribution [20, 32]. However, previous work has noted that MSM estimates of stationarity are biased by the trajectory training data [38, 4]. To partially address this issue, the MSM trajectories can be reweighted based on the matrix’s stationary solution [46]. This procedure — referred to here as “single-shot reweighting” — can be helpful. However, it cannot correct the weights of trajectories within the chosen discrete states, nor can it provide a set of weights consistent with a subsequently computed transition matrix [33, 34], potentially skewing the estimation of observables.
@@
Unbiased estimation only occurs in the traditional MSM framework when the trajectories in each discrete state are locally consistent with the stationary distribution for the chosen boundary conditions. For example, source-sink boundary conditions lead to the challenging requirement to sample from the resulting nonequilibrium steady state[33, 46, 11]. The need for unbiased equilibrium or nonequilibrium samples creates a chicken and egg problem, which can be addressed by an iterative solution [33, 34] at the cost of generating long trajectories. We note that non-traditional MSMs can be used to obtain unbiased estimates of observables [28], but traditional MSMs are biased for both equilibrium and nonequilibrium observables even when significant trajectory data is available [4, 42].
@@
As an approach for correcting standard MSMs, we here introduce the “Randomized ITErative trajectory reWeighting” (RiteWeight) algorithm. RiteWeight reweights trajectories generated without biasing forces into their correct stationary distribution when sufficient data is available. The method applies to both equilibrium and nonequilibrium steady states. It naturally uses trajectories of any length, as short as just a single time step, as it does not rely on dynamical relaxation. For example, RiteWeight can employ data generated from standard molecular dynamics or path sampling approaches, as long as no biasing forces have been used. We emphasize that RiteWeight is distinct from conventional importance sampling, which requires a known and well-sampled initial distribution [16, 50].
@@
Figure 1 shows how the RiteWeight algorithm iteratively reweights trajectories using the estimated stationary measure for a discrete-state transition matrix , which is determined in each iteration by a new random clustering. The relative weights of the trajectory segments in each cluster are fixed during a single iteration. In the equilibrium case, the transition matrix used is akin to a MSM based on weighted trajectories. In the nonequilibrium case, the matrix is computed from weighted trajectories conforming to source-sink boundary conditions, akin to a “history augmented” MSM [41, 40, 42, 11]. However, RiteWeight uses no history information and thus eliminates the requirement that all segments to be traced back to the source state. RiteWeight also does not require the Markov property to hold at the cluster level, because discrete states are used only to estimate stationarity and not to propagate dynamics over time.
@@
RiteWeight differs from prior strategies for reweighting sampled trajectory data. For instance, one approach explicitly approximates the configuration-space density function in order to reweight configurations [47]. Another method updates the weights using a variational principle for the equilibrium distribution [25]. RiteWeight, however, does not require density estimation and only uses standard Markov models without additional fitting parameters or assumed functional forms. In prior work, our research group has applied iterative reweighting without changing cluster definitions [34], but this approach does not yield the quasi-continuous distribution which RiteWeight achieves.
@@
The present report evaluates the RiteWeight algorithm through mathematical analysis and empirical testing. First, we present mathematical analysis identifying the fixed point of the RiteWeight algorithm. Then, we present numerical tests based on both synthetic molecular dynamics (SynMD) and true MD trajectories of the Trp-cage miniprotein. SynMD consists of trajectories generated from a fine-grained MSM with each state mapped to an atomistic configuration, and it enables comparison to an exactly calculable reference distribution [35]. The true MD data is a single 208 s trajectory generated by the Shaw group [26]. We compare the performance of RiteWeight, single-shot reweighting, and traditional MSMs for estimating the stationary distribution from mis-distributed data sets. Overall, RiteWeight yields better agreement with reference values for all observables considered in both equilibrium and nonequilibrium scenarios. RiteWeight achieves good performance using extremely short trajectory segments, enabling computation of mechanistic, path-based observables. Last, RiteWeight results are independent of the number of clusters used to discretize the configuration space.
@@ 2 RiteWeight Algorithm
The RiteWeight algorithm estimates steady state probability distributions using kinetic information extracted from numerous short MD trajectories or one long trajectory. The algorithm processes sequential pairs of configurations generated by unbiased dynamics and separated by a fixed lag time: . We refer to these pairs as “trajectory segments”. Unlike MSMs, the algorithm’s ability to estimate stationary probabilities is not constrained by the choice of lag time [32, 20]. Whereas single-shot reweighting requires within-cluster stationarity for unbiased results [33, 46, 11], RiteWeight imposes no such constraints. Therefore, transition pair data can be collected right from the start of the trajectories, ensuring that no data is discarded and all available information is leveraged in the analysis.
In the remainder of this report, we use a standardized nomenclature to avoid confusion. The discrete regions of configuration space employed by RiteWeight will be called “clusters,” while the discrete regions used by MSMs will simply be called “states.” Collections of states used in the analysis of nonequilibrium fluxes will be called “macrostates.” The term “microstate” will refer to finest resolution available for describing a system, such as points in configuration space. Thus, both clusters and states typically consist of multiple microstates, and a macrostate typically consists of multiple states.
@@ 2.1 The algorithm
Assign an initial weight to each trajectory segment connecting two consecutive configurations, using the normalization . A possible choice is uniform weighting: each where is the number of segments.
@@
Calculate the stationary probabilities for the clusters using the left leading eigenvector of the matrix , i.e., .
@@
(2) where is the previous weight of the trajectory segment, is the new weight, is the current iteration’s estimate of the stationary probability for the cluster , and . Also see the more general update rule in eq. (8), which introduces an additional learning rate parameter.
@@
Several points are noteworthy. In step 2, there are two approaches for assigning the initial weights to trajectory segments. One option is to assign uniform weights to all segments, which serves as an uninformative prior. Alternatively, if there is prior knowledge regarding the stationary distribution of the system, this information can be used to assign the initial weights in a more informed manner. Random clusters are defined in step 3, and a new clustering is introduced at every iteration. In step 6, the weight of each trajectory in cluster is multiplied by a fraction , where is the currently estimated stationary probability and is the previous estimate for the cluster.
Last, we define convergence in the numerical experiments by running RiteWeight until the weights or mean first passage time estimate exhibits small, steady fluctuations (typically – iterations). Then, we average the weights over the final 1,000 or 10,000 iterations to further reduce noise. See supplementary Figs. S1, S2, and S6 for examples of RiteWeight convergence analysis.
@@ 2.2 Analysis of RiteWeight fixed point
For simplicity and clarity, our analysis assumes that the configurations all belong to a discrete space of “microstates” indexed by or . We expect that RiteWeight in continuous space behaves similarly to RiteWeight in discrete space at a level of resolution automatically determined by the algorithm and its stopping criterion.
@@
Here, is the initial weight assigned to segment . The appendix proves the fixed point of RiteWeight equals the stationary vector of , assuming the stationary vector is unique and the clusters are chosen according to a random hyperplane tesselation.
The fixed point analysis suggests that RiteWeight will converge to the true stationary distribution given sufficiently dense and unbiased local sampling. In the limit of high transition counts, the trajectory segments will be described by a fine-grained transition matrix that approaches the true transition operator. Thus, RiteWeight will find the correct distribution regardless of the number of clusters or the lag time. However, in more realistic settings, the accuracy of RiteWeight is constrained by the data analyzed, pointing to a valuable role for the adaptive sampling methods that will be mentioned in the discussion section.
Given the description of the RiteWeight fixed point, it is natural to ask whether the RiteWeight stationary distribution can be derived directly from a fine-grained MSM transition matrix. In practice, this is not possible because the resolution of the microstates captured by RiteWeight cannot be easily determined for an arbitrary data set. The results section will show the limitations of simply using a fine-grained Markov state model without RiteWeight.
@@ 3 Test systems
Here we introduce the synthetic MD and atomistic MD systems used in our empirical RiteWeight experiments.
@@ 3.1 Synthetic MD for Trp-cage
Model. The SynMD model for the Trp-cage miniprotein [35] is derived from an MSM trained on a 208 s atomistic MD trajectory [26]. The MSM uses 10,500 states, which is a finer discretization than a typical MSM, and each state is mapped to a specific atomistic configuration. These 10,500 configurations are then used as microstates for SynMD. Kinetic Monte Carlo yields a “synthetic” MD trajectory that closely matches the statistical behavior of the original atomistic MD [35]. The MSM lag time and hence the interval between synMD configurations is 1 ns, and the trajectories can be processed using standard MSMs or RiteWeight. Because the synMD model is governed by an explicitly known MSM transition matrix, we can exactly calculate the equilibrium distribution for microstates using the leading left eigenvector of the matrix.
Trajectory preparation. We prepared a data set of 5 ns SynMD trajectories using a sampling measure far from equilibrium. First we applied tICA with a 5 ns lag time [42] to the set of minimal residue-residue distances, which are defined by the closest distances between the heavy atoms of two residues separated in sequence by at least two neighboring residues. Then we assigned indices to the 10,500 microstates based on the slowest time-lagged independent component (TIC1) [35]. We partitioned the SynMD indices into 500 bins and selected 20 initial microstates randomly with replacement from each bin. For each selected microstate, we initiated a short trajectory of 5 ns, i.e., 5 steps. Hence the unprocessed trajectory data consists of 10,000 trajectory segments, each 5 steps long, distributed roughly evenly along the tIC1 coordinate.
RiteWeight analysis. We applied RiteWeight in a manner blind to the discrete nature of synMD, using only the mapped atomistic configurations. We defined features using the slowest 10 tICs, and we rescaled the tICs based on commute distances. Then we performed RiteWeight using either or clusters based on the rescaled tICs.
@@ 3.2.1 Equilibrium distribution
Trajectory preparation. Again we prepared a data set of short trajectories using a sampling measure far from equilibrium. To that end, we applied tICA with a 10 ns lag time [42] to the set of minimal residue-residue distances. Then we projected all configurations from the 208 s trajectory onto tIC1 and grouped them into 100 uniformly spaced bins. We subsampled 10,000 configurations within each bin; however, if fewer than 10,000 configurations were available in a bin, we selected all the configurations. We used these configurations as the starting points for segments consisting of two snapshots from the long trajectory separated by a 10 ns lag time.
RiteWeight and MSM analysis. We retained a sufficient number of tICs to explain 95% of the total variance and rescaled the tICs based on commute distances. For RiteWeight, we performed random clustering () using the rescaled tICs as features. For the MSM analysis, we performed -means clustering based on – clusters. After clustering, we used the PyEMMA software package [39] for all MSM analysis.